OpenStudy (anonymous):

15n^2-n-28 factored with work

OpenStudy (amistre64):

n- /15 ; n+ /15 15(28) = 240+320+50 = 700

OpenStudy (amistre64):

1,700; 2,350; 4,125; 5,120, 10; 60, 20; 30 will work

OpenStudy (amistre64):

30 and 20 will not work; thats 10 not 1 lol n- 30/15 ; n+ 20/15 n- 2 ; n+ 4/3 n- 2 ; 3n+ 4

OpenStudy (amistre64):

14,50; 28,25..soo close

OpenStudy (amistre64):

n(n+1) = 700 n^2 +n -700 n= -1/5 +- sqrt(2801)/2

OpenStudy (anonymous):

so then what is it when factored out

OpenStudy (amistre64):

trying to ge to that, doesnt seem to work...did you miss a number like 3 perhaps?

OpenStudy (amistre64):

that middle terms looks to be missing a number to it

OpenStudy (anonymous):

no thats exactly how it is on my worksheet

OpenStudy (amistre64):

420...it helps if i know how to multiply lol..... back on track lol

OpenStudy (amistre64):

10,42; 20, 21 thats it lol

OpenStudy (amistre64):

n- 21/15 ; n+ 20/15 n- 7/5 ; n+ 4/3 5n- 7 ; 3n+ 4

OpenStudy (amistre64):

my georgia math got befuddled ;)

OpenStudy (anonymous):

lol that happens to me alot i was thinking for a second i would get some bonus points for bringing a mistake to the teachers attention lol

OpenStudy (amistre64):

15 28 ---- 240 160 50 ---- 450 ....im 30 off, why?

OpenStudy (anonymous):

hmm im not to sure why

OpenStudy (dumbcow):

20*15 = 300 8*10=80 8*5=40 sum =420